1.A 2 kg particle starts at the origin and moves along positive x axis.The net force acting on it measured at intervals of 1m are 28 30 45 48 50 64.5 13.5 12.2 40.8 27 and 23 newtons. Find the total work done during the whole process.
1
Expert's answer
2020-10-08T18:14:46-0400
"W=\sum (Fx)=\sum F \cdot Nd\\\sum F=28+0+40+48+50+64.5\\+13.5+12.2+40.8+27+23=347\ N"

"W=347(11)(1)=3817\ J"